# A Lean Dataset for International Math Olympiad: Small Steps towards Writing Math Proofs for Hard Problems

### v1

@@ 3.1 Prompting GPT-4
Here, we prompt GPT-4 with a statement requesting it to write a formal proof for the problem, instructing it to think step by step, explain the theorem, the approach, and the proof, following prompting techniques suggested by [15]. We also provide the formal Lean statement of the problems in the prompt so that issues about the formalization of the theorem statement will not arise. Furthermore, we crosscheck whether GPT-4 can explain the formal Lean statement in English, and whether the natural language description of the GPT-4 for each lemma was correct (see Figure G1). Therefore, providing the problem statement in Lean does not create confusion for the model as it can describe it correctly in English. Please see Appendix G to see the exact prompt structure and some examples.
Table 2 shows the evaluation of the responses of GPT-4 in natural language (NL) and in Lean. GPT-4 defaults/prefers to write formal proofs in Lean3 rather than Lean4. Hence, we evaluate its formal proofs directly in Lean3. The last column of this table evaluates whether there is a match between the proof written in NL and the one in Lean. GPT-4 may describe a proof in NL, but follow a different approach when writing the proof in Lean. For example, for the three lemmas corresponding to IMO 1959 P1 (the first row in Table 2), GPT-4’s proof in NL is correct for all the lemmas (100% in "Correct proof in NL"). However, GPT-4’s formal proof for one of these lemmas do not match the approach described in NL: 67% in "Match between NL and Lean". Indeed, GPT-4’s formal proofs for all these three lemmas have one or more mistakes: 0% in Correct proof in Lean.
Table 2: Analyzing the GPT-4’s responses to our lemmas
# Problem Correct proof in NL Correct proof in Lean Match between NL and Lean
1 1959-p1 100% 0% 67%
2 1960-p2 25% 25% 75%
3 1962-p2 73% 60% 60%
4 1964-p2 50% 40% 40%
5 1965-p2 36% 21% 15%
6 1983-p6 32% 13% 12%
7 1984-p6 6% 0% 5%
8 1992-p1 0% 0% -
9 1997-p5 0% 0% -
10 2022-p2 0% 0% -
11 2022-p5 0% 0% -
12 2023-p4 0% 0% -
@@
Moreover, Table 3 shows a more detailed evaluation of the formal proofs written by GPT-4. If a proof is completely correct and can pass the Lean system, it will be counted in the column "No error". Otherwise, a proof may have one or more of the following issues: hallucinations, wrong approach, wrong implementation, and minor errors.
@@
"Hallucination" refers to cases where the proof contains a lemmas or tactic that does not exist in the Mathlib library and it is also not defined separately in the proof. This happens more often when the model is uneducated on a topic. In such cases, GPT-4 usually assumes that there is one or a few magic lemmas in the Mathlib library that can prove the given lemma outright or in a few steps.
@@
"Wrong approach" refers to cases where the approach to prove a lemma is wrong. This means that either some of the individual proof steps are wrong or the steps do not logically lead to what the model wants to prove.
@@
"Wrong implementation" refers to cases where the approach itself is correct but the model does not use the correct procedures or it does not call the correct lemmas from the Mathlib library.
@@
Finally, "minor errors" refers to cases where the model uses a correct lemma or tactic from the Mathlib library but it does not provide the correct arguments to it, or it misses some of the arguments.
@@
Table 3: Analyzing and grading the errors in the Lean proofs written by GPT-4
@@
Hallucinations Wrong approach Wrong implementation Minor errors
1 1959-p1 0% 67% 33% 67% 67%
2 1960-p2 12% 62% 75% 88% 88%
3 1962-p2 33% 40% 40% 47% 67%
4 1964-p2 10% 40% 50% 60% 90%
5 1965-p2 11% 71% 65% 79% 94%
6 1983-p6 0% 84% 91% 100% 100%
7 1984-p6 0% 100% 100% 100% 100%
8 1992-p1 0% 100% 100% 100% 100%
9 1997-p5 0% 100% 100% 100% 100%
10 2022-p2 0% 100% 100% 100% 100%
11 2022-p5 0% 100% 100% 100% 100%
12 2023-p4 0% 100% 100% 100% 100%
3.2 Interacting with GPT-4: Hinting and Chain of Thoughts
We then try to improve the response of GPT-4 by interacting with it. For each incorrect proof, we provide feedback, pointing out the mistakes, and asking the model to fix the issues. If the approach is wrong, we mention that the approach is wrong. If there are hallucinations, we point out such mistakes as well. The results are presented in Table 4.
Table 4: Interacting with and hinting GPT-4 to improve its proof
1 1959-p1 0% 67% 67%
2 1960-p2 12% 25% 25%
3 1962-p2 33% 47% 47%
4 1964-p2 10% 30% 30%
5 1965-p2 11% 17% 17%
6 1983-p6 0% 7% 8%
7 1984-p6 0% 8% 8%
8 1992-p1 0% 0% 0%
9 1997-p5 0% 0% 0%
10 2022-p2 0% 1% 1%
11 2022-p5 0% 0% 0%
12 2023-p4 0% 0% 0%
We interact with GPT-4 at most 10 times. If the model does not succeed in writing a correct proof after 10 rounds of feedback, we stop the process. There are papers in the literature that go as far as hinting the GPT 100 times. We find such an approach labor-intensive and not scalable. It is worth noting that 10 times can already be costly without bringing fruits. Moreover, we have observed that within that range of 10 times, GPT-4 has made no significant improvements.
Later in this paper, in the section on source criticism, we show that most of the GPT-4’s correct answers are already available on the Internet. Hence, a search engine that can retrieve from the contents of the Mathlib library or simply the contents on the Internet may be more efficient than interacting with GPT-4 many times over.
The results further suggest that GPT-4 can be considered uneducated on many of the subjects listed in Table C1, especially the topics related to IMO problems between 1984 to 2023. One might imagine that having LLMs trained on these topics may have the potential to address some of the shortcomings. However, since the search space and the number of combinations of lemmas in topics such as number theory are immensely large, unlike the problems in geometry, it remains a questions whether the best model to write formal proofs for problems in number theory and algebra will be an LLM or some other method that primarily relies on searching the existing lemmas in the Mathlib library.
We also perform manual proof repair as described in Appendix I.
4 Case study: Insights about the dataset
To provide a better understanding of the dataset and GPT-4’s abilities, in this section, we provide a case study on a small portion of the proof from the IMO 1997 P5. Note that in this section, we follow a chain of simplifications of the problem for the purpose of demonstration. Unlike this section, the results in section 3.2 follow a specific protocol of asking GPT-4 to prove one specific lemma, and repeating the same request while providing feedback about the mistakes of the previous response.
Out of the 390 lines of Lean proof for this problem, we zoom in on only 80 lines of the proof. Here is the lemma:
@@ 4 Case study: Insights about the dataset
In the first attempt, GPT-4 gives a very long answer, reviewing all the items in the lemma, but does not provide any proof, not in natural language and not in Lean. It does not know how to approach the problem.
For the proof, one needs to first show that $2y^{2}<x$. After proving this inequality, we can rely on prime factorization of both sides of the inequality to prove the target of the lemma. To guide the GPT-4, we directly ask if it can prove $2y^{2}<x$. Again, it goes on reviewing the statements line by line, providing a large amount of explanations, but, eventually, does not find a way to prove it – again no proof in Lean and no proof in natural language.
We then go back to the first statement and take it for granted that $2y^{2}<x$. This time, GPT-4 attempts to provide a proof, a simple proof relying on a lemma called nat.dvd_of_factorization_eq. Such lemma does not exist in the Mathlib library so its use can be considered a hallucination, perhaps out of desperation. We point out this issue to GPT-4, which leads it to write another proof including this:
rw nat.factorization_mul_of_coprime (nat.coprime.symm (nat.prime.coprime_pow_two_of_ne (ne_of_lt hyx) (nat.prime_of_mem_factorization (exists_mem_of_ne_empty (factorization_pos_of_pos y_pos)))) x_pos) at $h_{2}$,
This line contains many non-existent lemmas. We ask GPT-4 to explain the lemmas one by one which it does happily not realizing that they are hallucinations. As an example, here is the description of GPT-4 for the non-existent lemma nat.prime.coprime_pow_two_of_ne:
"This lemma asserts that if $a\neq b$, and a is a prime, then $a$ is co-prime to $b^{2}$. This is based on the property that a prime number does not divide any power of another distinct number."
GPT-4’s explanation is completely wrong and implies a lack of understanding about the topic.
We further simplify the problem by granting the statement that
$x\bullet 2\bullet nat.factorization\;y\leq x\bullet nat.factorization\;x.$
This statement by itself needs several lines of proof, but we grant it to GPT-4 to make the problem easier. GPT-4 again goes back to its original hallucination of the lemma nat.dvd_of_factorization_eq. We inquire about this lemma and GPT-4 explains that if two numbers have equal prime factorizations, then those numbers will divide each other. We ask GPT-4 if it has been able to establish that $y^{2}$ and $x$ are equal. It answers no and apologizes. Then we point out that nat.dvd_of_factorization_eq does not even exist. It apologizes again.
Finally, to step aside from prime factorization, we simplify the problem further by adding a new statement $(y^{2})^{x}|x^{x}$, and ask GPT-4 to prove $y^{2}|x$. This time, it ignores the new statement, provides a wrong proof in natural language, and does not write any proof in Lean.
What we can conclude from this case study is that GPT-4 does not have adequate knowledge about this topic, neither in natural language, nor in Lean. Anytime it tries to come up with an answer, it hallucinates lemmas that can prove the problem outright, but those lemmas do not exist, and at the same time, they are incorrect or inapplicable. So, it is not viable to prove those lemmas and then use the rest of the GPT-4’s proof.
This case study is insightful, not as a failure case of GPT-4, but as a showcase of the gap between the GPT-4’s abilities and what it takes to write the formal proof for the IMO 1997 P5. This lemma was just a small portion of the proof for this IMO problem, and until we have a model that can prove this lemma, we will not be able to write the complete proof for this IMO problem.

### v2

@@ 3 Evaluating the ability of LLMs in proving our lemmas
We prompt o3-mini with a statement requesting it to write a formal proof for the problem, instructing it to think step by step, explain the theorem, the approach, and the proof, following prompting techniques suggested by Kojima et al. (2022). We also provide the formal Lean statement of the problems in the prompt so that issues about the formalization of the theorem statement will not arise. Furthermore, we verify that o3-mini’s explanation of the problems in natural language matches the provided Lean statements. Table 2 shows the evaluation of the responses of o3-mini in natural language (NL) and in Lean. The last column of this table evaluates whether there is a match between the proof written in NL and the one in Lean. A LLM may describe a proof in NL, but follow a different approach when writing the proof in Lean.
Table 2: Analyzing the o3-mini’s responses to our lemmas
# Problem # of lemmas Correct proof in NL Correct proof in Lean Match between NL and Lean
1 1959-p1 4 100% 50.0% 100%
2 1960-p2 9 55.6% 11.1% 100%
3 1962-p2 14 92.9% 42.9% 100%
4 1964-p2 9 77.8% 33.3% 100%
5 1965-p2 73 97.3% 16.4% 100%
6 1983-p6 53 64.2% 34.0% 100%
7 1984-p6 64 73.4% 20.3% 100%
8 1985-p6 427 75.2% 19.7% 95.6%
9 1992-p1 91 87.6% 27.5% 100%
10 1997-p5 122 69.7% 24.6% 100%
11 2022-p2 61 77.0% 41.0% 100%
12 2022-p5 265 63.4% 22.6% 92.8%
13 2023-p4 137 88.3% 27.0% 92.7%
total 1,329 75.5% 23.8% 96.4%
We provide up to 10 rounds of feedback to o3-mini in the instances that it does not write a correct proof. The feedback is automatically generated using REPL [^1] https://github.com/leanprover-community/repl, a library in Lean which summarizes the errors when compiling a Lean file. The results are presented in Table 3 demonstrating the effectiveness of our automated feedback system in improving the accuracy of o3-mini.
Table 3: The effect of automatic feedback loop on the accuracy of o3-mini
1 1959-p1 50.0% 50.0% 50.0%
2 1960-p2 0.0% 11.1% 11.1%
3 1962-p2 7.1% 21.4% 42.9%
4 1964-p2 22.2% 33.3% 33.3%
5 1965-p2 0.0% 8.2% 16.4%
6 1983-p6 9.4% 20.8% 34.0%
7 1984-p6 3.1% 14.1% 20.3%
8 1985-p6 4.9% 16.2% 19.7%
9 1992-p1 5.5% 16.5% 27.5%
10 1997-p5 4.9% 16.4% 24.6%
11 2022-p2 16.4% 36.1% 41.0%
12 2022-p5 4.5% 18.1% 22.6%
13 2023-p4 10.2% 24.8% 27.0%
total 6.0% 18.4% 23.8%
@@
Moreover, Table 4 shows a detailed evaluation of the formal proofs written by o3-mini at its final attempt. If a proof is completely correct and can pass the Lean system, it will be counted in the column “No error”. Otherwise, a proof may have one or more of the following issues: hallucinations, wrong approach, wrong implementation, incomplete proof, and/or minor errors.
@@
“Hallucination” refers to cases where the proof contains one or more lemmas or tactics that do not exist in the Mathlib library and it is also not defined separately in the proof. This happens more often when the model is uneducated on a topic. In such cases, LLM may assume that there is one or a few magic lemmas in the Mathlib library that can prove the given lemma outright or in a few steps.
@@
“Wrong approach” refers to cases where the approach to prove a lemma is wrong. This means that either some of the individual proof steps are wrong or the steps do not logically lead to what the model wants to prove.
@@
“Wrong implementation” refers to cases where the model does not use the correct procedures or it does not call the correct lemmas from the Mathlib library.
“Incomplete proof” refers to cases where the proof is explicitly or implicitly incomplete. An explicitly incomplete proof contains one or more sorry statements, whereas an implicitly incomplete proof does not contain such statements but is unfinished.
@@
Finally, “minor error” refers to cases where the model uses a correct lemma or tactic from the Mathlib library but it does not provide the correct arguments to it, or it misses some of the arguments.
@@
Table 4: Analyzing the errors in the Lean proofs written by o3-mini
@@
Hallucination Wrong approach Wrong implementation Incomplete proof Minor error
1 1959-p1 50.0% 50.0% 0.0% 50.0% 0.0% 50.0%
2 1960-p2 11.1% 33.3% 44.4% 11.1% 0.0% 88.9%
3 1962-p2 42.9% 7.1% 7.1% 28.6% 0.0% 57.1%
4 1964-p2 33.3% 22.2% 22.2% 0.0% 11.1% 44.4%
5 1965-p2 16.4% 6.8% 2.7% 26.0% 1.4% 83.6%
6 1983-p6 34.0% 0.0% 35.8% 35.8% 3.8% 60.4%
7 1984-p6 20.3% 0.0% 26.6% 20.3% 3.1% 79.7%
8 1985-p6 19.7% 0.5% 18.0% 53.4% 14.8% 79.9%
9 1992-p1 27.5% 1.1% 12.1% 48.4% 1.1% 72.5%
10 1997-p5 24.6% 4.1% 30.3% 31.1% 6.6% 75.4%
11 2022-p2 41.0% 6.6% 23.0% 23.0% 0.0% 59.0%
12 2022-p5 22.6% 2.3% 32.1% 58.1% 9.4% 74.3%
13 2023-p4 27.0% 0.0% 1.5% 63.5% 12.4% 73.0%
total 23.8% 2.3% 20.4% 46.9% 9.0% 75.1%
3.2 Performance of SOTA models on the dataset
We then evaluate the state-of-the-arts models in the literature. We consider Goedel-Prover as it has the highest accuracy on the miniF2F test set. We also consider DeepSeek Prover 1.5 which is the base of the Goedel-Prover and it used to be the model with the highest accuracy before the release of Goedel-Prover. Finally, we consider ReProver, with and without retrieval, which is an established model in the literature with an open-source training set. We note that both Goedel-Prover and DeepSeek Prover do not have an open-source training set while ReProver has. Table 5 shows the accuracy of these models on our lemmas based on the IMO problem where the lemmas are extracted from.
Table 5: Comparing performance of SOTA models on our lemmas dataset (blue shows the IMO problem that a model has the best accuracy on its extracted lemmas, and red marks the IMO problem that a model has the worst accuracy on its extracted lemmas. underline marks the model with the highest accuracy on the lemmas extracted from each IMO problem.)
Problem # of lemmas DeepSeek Prover-v1.5-RL (@32) Goedel-Prover (@32) ReProver retrieval ✗ ReProver retrieval ✓ o3-mini (with 10 e.f.)
1959-p1 4 3 (75.0%) 2 (50.0%) 2 (50.0%) 2 (50.0%) 2 (50.0%)
1960-p2 9 7 (77.8%) 6 (66.7%) 3 (33.3%) 4 (44.4%) 1 (11.1%)
1962-p2 14 13 (92.9%) 12 (85.7%) 7 (50.0%) 8 (57.1%) 6 (42.9%)
1964-p2 9 5 (55.6%) 5 (55.6%) 5 (55.6%) 5 (55.6%) 3 (33.3%)
1965-p2 73 48 (65.8%) 47 (64.4%) 47 (64.4%) 46 (63.0%) 12 (16.4%)
1983-p6 53 25 (47.2%) 32 (60.4%) 28 (52.8%) 29 (54.7%) 18 (34.0%)
1984-p6 64 31 (50.0%) 33 (51.6%) 25 (39.1%) 24 (37.5%) 13 (20.3%)
1985-p6 427 116 (27.2%) 116 (27.2%) 89 (20.8%) 89 (20.8%) 84 (19.7%)
1992-p1 91 48 (52.7%) 54 (59.3%) 35 (38.5%) 34 (37.4%) 25 (27.5%)
1997-p5 122 51 (41.8%) 49 (40.2%) 48 (39.3%) 51 (41.8%) 30 (24.6%)
2022-p2 61 34 (55.7%) 30 (49.2%) 24 (39.3%) 25 (41.0%) 25 (41.0%)
2022-p5 265 89 (33.6%) 76 (28.7%) 80 (30.2%) 77 (29.1%) 60 (22.6%)
2023-p4 137 52 (38.0%) 41 (29.9%) 43 (31.4%) 45 (32.8%) 37 (27.0%)
Total 1,329 522 (39.3%) 504 (37.9%) 436 (32.8%) 439 (33.0%) 316 (23.8 %)
The model with the best accuracy on our lemmas turns out to be the DeepSeek-Prover with accuracy of 39.3%. Although Goedel-Prover has a higher accuracy on miniF2F test set, it shows a lower performance on our dataset, proving 18 lemmas fewer than DeepSeek. At the same time, ReProver with retrieval proves 83 lemmas fewer than DeepSeek corresponding to accuracy of 33%.
o3-mini proves a total of 316 lemmas corresponding to accuracy of 23.8%. This accuracy is lower than the accuracy of the other three LLMs, yet it seems impressive given that, unlike the previous models, o3-mini is a general purpose model, not specifically designed for proving mathematical theorems in Lean language. This indicates a high potential for o3-mini to achieve SOTA accuracy on our dataset and on miniF2F after being fine-tuned on appropriate data such as the one used for fine-tuning Goedel Prover.
Additionally, we evaluated the accuracy of Qwen Math 2.5 (Yang et al., 2024) and Llama 3.1 (Dubey et al., 2024) on our lemmas in formal language. Qwen Math 2.5 was able to prove only 1 lemma while Llama failed to prove any of the lemmas. We note that these two models are not designed specifically for proving theorems in formal language, nevertheless, their ability for automated theorem proving significantly lag behind o3-mini. The complete results are shown in Table E4.
3.3 Analyzing the proof lengths
Here, we look at the patterns in the correct proofs written by the LLMs, specifically with regard to the proof lengths. Figure 1 shows the distribution of the length of correct proofs written by LLMs for the lemmas in our dataset, and Table 6 shows the accuracy of each LLM in proving the subsets of lemmas for various ranges of proof lengths. Out of the 1,329 lemmas in our dataset, Goedel Prover can prove 504 lemmas correctly. What is common among these 504 lemmas is that they all have short proofs, majority of them 1 liners using the automatic Lean solvers such as nlinarith. The longest correct proof that Goedel Prover has written for our lemmas has 15 lines and the longest proof written by DeepSeek-Prover has 12 lines. Surprisingly, the longest correct proof is written by o3-mini consisting of 34 lines.
For the lemmas that Goedel Prover and DeepSeek Prover can correctly prove, the average length of the proof is less than 3 lines, and the maximum length of 15 and 12 lines are several standard deviations away from that mean (See Appendix F for further details). Hence, writing a correct proof of 15 or 12 lines is more of an anomaly for Goedel and DeepSeek. Overall, one can conclude that the ability of these two LLMs in writing correct proofs seems to be limited only to short proofs. These observations suggest a low capability for planning and reasoning in these LLMs as writing a long proof requires devising an elaborate plan as a prerequisite.
On the other hand, o3-mini shows a higher capability in writing correct long proofs in Lean as shown in Figure 1. While o3-mini fails to write the correct proof for most of the 1-liners, it does considerably better than Goedel and DeepSeek in writing sensible long proofs in Lean. The average length of the correct proofs that o3-mini writes is 7.6 lines. Note that in Table 6, the proof lengths are based on the proofs in our dataset which tend to be shorter than the proofs written by LLMs. Based on those lengths, the accuracy of Goedel and DeepSeek is 0% for all the lemmas with proofs longer than 10 lines. o3-mini is the only model that proves some of such lemmas, achieving an accuracy of 5.8% for lemmas with proof length of 11 to 15 and an accuracy of 1.9% for lemmas with proof length of 26 to 100 lines. For lemmas with proof length of 6 to 11 lines, the best accuracy is 46% by Goedel, and for lemmas with proof length of 6 to 10 lines, the best accuracy is 10% by o3-mini. The only type of lemmas that DeepSeek performs better are the lemmas that have a proof of 1 or 2 lines. Even for those lemmas with very short proofs, none of the models can achieve a near perfect accuracy. All these observations hints at opportunities for improving these models.
Figure 1: Distribution of the length of correct proofs written by LLMs. The histogram also includes the distribution of human-written proof lengths for comparison.
Table 6: Accuracy of models on our lemmas based on the length of proofs
range of proof length 1-2 3-5 6-10 11-15 16-25 26-100 101-298
# of lemmas 497 204 169 103 128 210 18
Deepseek-Prover-v1.5-RL 84.3% 42.6% 8.9% 0.0% 0.0% 0.0% 0.0%
Goedel-Prover 79.5% 46.1% 8.9% 0.0% 0.0% 0.0% 0.0%
ReProver retrieval ✗ 81.7% 11.8% 0.0% 0.0% 0.0% 0.0% 0.0%
ReProver retrieval ✓ 82.9% 9.8% 0.0% 0.0% 0.0% 0.0% 0.0%
o3-mini 49.7% 24.0% 10.1% 5.8% 0.0% 1.9% 0.0%
We also observe that many of the correct proofs written by Goedel-Prover and DeepSeek-Prover contain a considerable number of irrelevant lines that can be purged from the proofs. Appendix G shows two examples of such proofs where large number of lines are used to prove irrelevant hypothesis or to repeat what is already proved or given. We refer to those proofs as bloated, and such irrelevant lines are not counted in the numbers reported in Figure 1 and Appendix F.
4 Decomposing a proof into its building blocks and extracting new lemmas
Here, we explain the procedures for lemma decomposition, and also derive upper bounds about the number of lemmas that can be extracted from an existing proof. The lemmas in our dataset were created by breaking down the proofs of the IMO problems. These proofs usually have specific structures that we exploit when breaking them down. First, we explain, how we exploit the structure of the proofs. Later, we will consider an extreme case and derive an upper bound on the number of lemmas that can be extracted from an $n$ line proof.
4.1 Breaking down the structure of a proof
Let us consider part of the proof for IMO 1997 P5 as an example. Here is the lemma:
@@ 4.1 Breaking down the structure of a proof
First, we can consider each of the intermediate hypotheses and define them as a separate lemma. The intermediate hypotheses in the proof are:
  • Lemma 1: prove $y^{2}<x$
  • Lemma 2: prove $2*y^{2}<x$
  • Lemma 4: $(y^{2})^{x}$ divides $x^{x}$
Each of these steps can be taken as a lemma.
In an extreme case, a proof may have $k=(n-1)/3$ intermediate hypothesis (each of them with 1 line of statement and 2 lines of proof leading to the 3 in the denominator) which would lead to $2/3*(n-1)$ extracted lemmas. For example, in this setting, a proof with 4 intermediate hypothesis may have $4*3+1=13$ lines. And that leads to $2/3*(13-1)=8$ extracted lemmas.
4.2 Case of a proof with no intermediate structure
The forward path would start with the first line of the proof, apply the line, obtain the new proof state, and define that as a new lemma. The proof for this new lemma is the $(n-1)$ lines of the proof that remain. Continuing this way, the forward path can extract $(n-2)$ lemmas where each lemma has at least 2 lines of proof. Here, a minimum 2 lines of proof can be used as a threshold on the easiness of lemmas.